# Concept Unlearning by Modeling Key Steps of Diffusion Process

### v2

@@ 4.3 NSFW Unlearning
As shown in Table 2, KSCU balanced unlearning effectiveness and generative retainability, achieving a UA of 96.5%, surpassing other methods in forgetting capability. Meanwhile, its FID-30K of 14.1 was the lowest among all baselines, confirming that KSCU best generative retainability. In contrast, although Salun achieves the highest UA, its FID of 67.9 reveals a significant degradation in generative quality. Moreover, we observed extensive large-area distortions in Salun’s generated results, indicating severe over-unlearning. Compared to methods with similar FID-30K scores, such as ESD and CA, KSCU demonstrated superior forgetting performance. Conversely, while methods like EDiff and Salun matched KSCU in unlearning effectiveness, they fell short in retaining generative ability. These results established KSCU as the most effective and balanced approach for NSFW concept unlearning.
@@
Robustness. In the adversarial sample defense experiment, excluding Salun (which exhibited over-unlearning), KSCU achieved the lowest ASR against P4d [6] and UDA [58] among all other SOTAs, demonstrating its strong robustness. On the other hand, we observed a certain degree of correlation between the performance of different methods under adversarial attacks and their forgetting precision. We believed this was because the i2p prompts used in the experiment were manually selected, diverse prompts with inherent adversarial properties. As a result, the unlearning performance on i2p can, to some extent, reflect the robustness of a given method in complex real-world scenarios or against adversarial samples.
4.4 Instance Unlearning
Figure 6: KSCU vs. ESD in Instance Unlearning. KSCU achieves forgetting with only 20% of ESD’s iterations. Despite ESD’s greater adjustments, KSCU demonstrates superior performance in erasing instance information.
To demonstrate KSCU’s effectiveness in unlearning instances memorized by diffusion models, we qualitatively compared the unlearning performance of KSCU and ESD through visualized instance unlearning results. The reason for choosing ESD was that the architecture of ESD was more similar to ours and ESD had shown good performance in previous experiments.
As shown in Figure 6, KSCU demonstrated superior performance in unlearning public figures. In KSCU’s results, facial features of public figures were effectively altered, whereas ESD partially retained the original characteristics. For artificial objects, ESD’s unlearning effect was barely noticeable, whereas KSCU replaced the target with either a visually similar but semantically distinct object (e.g., Eiffel Tower $\rightarrow$ highway) or a semantically related alternative (e.g., Titanic $\rightarrow$ small boat). KSCU, due to its well-designed approach, achieved a more effective instance unlearning performance. More importantly, KSCU achieved this with lower computational cost, highlighting its potential for broader applications.
4.5 Ablation Study
@@
Table 3: Ablation Study: Impact of Prompt Augmentation(PA) and Key Step Table (with Different Start $S$) on KSCU’s Performance
Method UA(%$\uparrow$ P4D(%$\downarrow$ UDA(%$\downarrow$ FID-30$\downarrow$
ESD [12] 88.2 39.3 73.7 15.4
KSCU w/o PA 92.5 30.3 51.8 14.5
KSCU $S$=20 79.7 59.2 85.3 15.7
KSCU $S$=10 96.1 23.0 48.1 14.3
KSCU $S$=0 96.0 26.7 54.5 18.8
KSCU 96.5 23.2 47.2 14.1
In the ablation study, we mainly evaluated the impact of the Key Step Table and Prompt Augmentation. After removing Prompt Augmentation, KSCU exhibited a decrease in UA and a significant increase in ASR against adversarial examples, indicating a reduction in model robustness. This strongly confirmed that Prompt Augmentation enhanced the model’s ability to better perceive the semantic space of the target concept, achieve precise concept erasure, and improve robustness.
In the ablation experiment of the Key Step Table’s start, we observed that when $S$=10 (covering 80% of steps), the performance was similar to $S$=15 (covering 70% of steps). However, when $S$=20 (covering 60% of steps), the unlearning effect of KSCU significantly decreased, suggesting that steps beyond 15 played a key role in the imaging of “nudity.” In the $S$=0 (covering 100% of steps) experiment, the unlearning performance remained strong. However, since all steps were affected, the FID increased, indicating a more significant negative impact on the model’s generative ability. Notably, the unlearning effectiveness of $S$=0 (100% steps), $S$=10 (80% steps), and $S$=15 (KSCU) was highly similar. This suggested that the first 30% of steps had little influence on the final rendering of ”nudity,” making their unlearning unnecessary. The results of the ablation experiment further demonstrated that for concept unlearning, KSCU, which focuses on unlearning key steps, was superior to the previous method of adjusting parameters for all steps.
5 Conclusion
This study investigates the impact of denoising steps on concept unlearning in diffusion models and proposes KSCU, a method that targets key denoising steps for concept unlearning. Through experiments on four different concept unlearning tasks, we demonstrate that KSCU surpasses previous state-of-the-art methods in terms of unlearning effectiveness, robustness, and generative retainability. The core ideas behind the key steps introduced in this work provide valuable insights for research in related fields.
Appendix
@@ Appendix A Analysis of Frequency Component Destruction in Diffusion Models
A.1 Definition of the Diffusion Process
@@
$d\mathbf{x}_{t}=g(t)d\mathbf{w}_{t},\quad\mathbf{x}_{0}\text{ as the initial condition.}$. (9)
A.2 Frequency-Domain Analysis
@@
A.3 Derivation of Frequency Component Degradation
A.3.1 Representation of the Initial Signal and Noise
@@
A.3.2 Power Spectral Density of the Signal
@@
[…]
A.3.3 Signal-to-Noise Ratio (SNR) Analysis
@@
A.3.4 Frequency Dependence of SNR Decay
@@
[…]
@@
A.4 Conclusion
The analysis confirms that, in the forward diffusion process, high-frequency components are degraded first, followed by lower-frequency components. Consequently, in the reverse denoising process, diffusion models reconstruct low-frequency structures first before refining high-frequency details. This frequency-dependent degradation provides theoretical justification for the progressive nature of generative sampling in diffusion models.
@@ Appendix B More Experimental Details
Appendix C Target Concept Unlearning
In Section 3.2, we derived our loss function, which ensures that the model samples in the direction opposite to the original concept $c$. This approach effectively prevents the generation of $c$ by replacing it with an unrelated concept. However, the replacement concept is uncontrolled, leading to unpredictable visual outputs. To address this, we introduce a slight modification to the loss function, guiding the model to replace $c$ with a specified target concept $c^{-}$, rather than an arbitrary alternative. The modified formulation is as follows:
$\begin{split}\hat{\epsilon}^{*}_{\theta}(\mathbf{x}_{t},t,\mathbf{c})&=\epsilon_{\theta}(\mathbf{x}_{t},t)+w\cdot\underbrace{(\epsilon_{\theta}(\mathbf{x}_{t},t,\mathbf{c^{-}})-\epsilon_{\theta}(\mathbf{x}_{t},t))}_{\text{Guiding towards }c^{-}}\\ &=\epsilon_{\theta}(\mathbf{x}_{t},t)+w\cdot\underbrace{(\epsilon^{*}_{\theta}(\mathbf{x}_{t},t,\mathbf{c^{+}})-\epsilon_{\theta}(\mathbf{x}_{t},t))}_{\text{Guiding towards }c^{+}}\end{split}$ (18)
where $c^{+}$ represents the concept to be unlearned, and $c^{-}$ denotes the target replacement concept. By applying the following modified loss function, we explicitly enforce the substitution of $c^{+}$ with $c^{-}$:
This formulation allows precise control over the replacement concept, ensuring that the model not only forgets $c^{+}$ but also consistently generates the desired substitute $c^{-}$, thereby enabling targeted concept substitution in diffusion models.
In the Figure 8, we present the results of KSCU’s target concept unlearning.
Appendix D Multi-concept Unlearning
To evaluate the effectiveness of KSCU in multi-concept unlearning, we provide visualized results 7. Our observations indicate that KSCU maintains strong performance in removing multiple concepts.
@@ Appendix D Multi-concept Unlearning
Appendix E More Visualization Results
Figure 8: Visualization results of KSCU with target replacement. The left side is SDv1.5, and the right side is the result of KSCU. The target concepts are bird and Van Gogh, and the replacement concepts are balloon and sketch.
Figure 9: KSCU’s visualization results for “nudity”, with SDv2.1 on the left and KSCU on the right.
Figure 10: visualization results of KSCU about artist styles, SDv2.1 on the left and KSCU on the right.
Figure 11: Visualization results of artist style erasing, model is SDv1.5

### v3

@@ IV-C Nudity Unlearning & Robustness
As shown in Table III, KSCU achieves a overall score of 0.92 in the nudity unlearning task, significantly outperforming all other methods. This indicates a strong balance between unlearning effectiveness and generative retainability. Specifically, KSCU attains a UA of 96.5%, surpassing most baselines in forgetting capability. At the same time, its FID-30K reaches 14.1—the lowest among all compared methods—demonstrating KSCU’s superior generative retainability. In contrast, although Salun and ANT achieve the highest UA, its FID of 67.9 reveals a significant degradation in generative quality. Moreover, we observed extensive large-area distortions in their generated results, indicating severe over-unlearning. Compared to methods with similar FID-30K scores, such as ESD and RECE, KSCU demonstrated superior forgetting performance. Conversely, while methods like EDiff and Salun matched KSCU in unlearning effectiveness, they fell short in retaining generative ability. These results established KSCU as the most effective and balanced approach for nudity unlearning, and also the fastest among all fine-tuning-based methods in terms of training efficiency.
@@
Robustness. In the adversarial sample defense experiment, excluding Salun and ANT (which exhibited over-unlearning), KSCU achieved the lowest ASR against UDA [54] among all other SOTAs, demonstrating its strong robustness. We also observed a certain degree of correlation between the performance of different methods under adversarial attacks and their forgetting precision. We believed this was because the i2p prompts used in the experiment were selected, diverse prompts with inherent adversarial properties. As a result, the unlearning performance on i2p can, to some extent, reflect the robustness of a given method in complex real-world scenarios or against adversarial samples.
IV-D Instance Unlearning
To further demonstrate the effectiveness of KSCU in unlearning specific instances memorized by diffusion models, we qualitatively compared KSCU with several state-of-the-art approaches using visualized instance unlearning results. The comparison provides a more holistic view of the trade-off between forgetting accuracy and generative retainability. The experiments reveal that, compared to other methods, KSCU not only removes the target instance more effectively but also achieves superior preservation of unrelated generative capabilities, thereby striking a better balance between forgetting and retention.
As illustrated in Figure 6, KSCU consistently excels at unlearning public figures. In KSCU’s outputs, the background, lighting, and body posture remain largely unchanged, while the facial features of the public figure are effectively altered or replaced, thus achieving successful instance unlearning without degrading overall scene integrity. By contrast, RECE is able to suppress the identity of the public figure but does so at the cost of severely disrupting unrelated visual content, leading to unnatural or distorted backgrounds. DUO and ANT provide only limited suppression of the target instance, often leaving recognizable traces of the original identity, while simultaneously introducing distributional shifts that reduce the fidelity and stylistic consistency of the generated image. ESD, on the other hand, shows relatively strong preservation of unrelated content but fails to sufficiently erase instance-level information, resulting in outputs that still closely resemble the original subject.
The comparison is even more striking in the case of landmark unlearning. DUO and ESD exhibit negligible forgetting effects, essentially leaving the landmark unchanged. RECE aggressively removes the landmark but does so in an indiscriminate manner, often erasing both the target concept and unrelated contextual elements, which substantially compromises realism. ANT is able to achieve more noticeable unlearning, but at the cost of visual artifacts and degraded generation quality, reflecting a lack of fine-grained control over what is erased. In contrast, KSCU demonstrates a more controlled behavior: it effectively suppresses the memorized landmark while replacing it with visually plausible alternatives, such as visually similar but semantically distinct objects (e.g., Eiffel Tower $\rightarrow$ highway) or semantically related but non-identical substitutes (e.g., Titanic $\rightarrow$ small boat). Such substitutions ensure that the generated images remain coherent and visually compelling, even after the target instance is forgotten. Among all evaluated methods, KSCU achieves the most effective and stable unlearning, while simultaneously producing the highest-quality generations.
These observations highlight that KSCU not only suppresses memorized instances but also substitutes them with plausible alternatives, thereby preserving overall image realism. More importantly, KSCU achieves instance unlearning with only 200 parameter updates, and its substantially reduced iterations and lower computational overhead underscore its potential scalability to large-scale unlearning tasks.
IV-E Ablation Study
In our ablation study, we primarily evaluated the three major components of KSCU: the Key Step Table, Prompt Augmentation, and Key Step Unlearning Optimization.
For the Key Step Table ablation, we varied the $S$ and $E$ indices in Algorithm 1, adjusting the step length $L$ to maintain comparable training frequencies. This design ensures that performance differences arise primarily from changes in step coverage, rather than training budget. By comparing KSCU variants with identical step coverage but different emphasis—such as early 70% vs. last 70%, early 80% vs. last 80%, and shrink $E$ vs. increase $s_{\text{cur}}$ (Table IV, Line 10)—we observe that configurations emphasizing later steps consistently achieve higher unlearning accuracy and robustness. This is intuitive: later denoising steps in diffusion sampling capture fine-grained semantic details that directly determine the fidelity of concept representation. Hence, unlearning at later steps is inherently more effective.
As shown in Table IV, configurations covering 100% of steps, the last 80%, and the last 70% achieve comparable unlearning effectiveness. However, unlearning accuracy drops by 16.8% when only the last 60% of steps are used. These results lead to two key insights: (1) the earliest 30% of steps contribute minimally to the final concept depiction (e.g., nudity) and can be excluded with little impact, thereby saving computation; and (2) steps after the 15th are critical for encoding the target concept and play a dominant role in unlearning effectiveness. Another notable observation is that although full-step unlearning achieves slightly higher accuracy, it incurs a larger FID due to broader interference with the model’s generative capacity. These findings highlight the efficiency and lower destructiveness of KSCU’s targeted fine-tuning compared to conventional approaches that indiscriminately modify parameters across all denoising steps.
@@ IV-E Ablation Study
TABLE IV: Ablation Study: Impact of Key Step Table (with different $S$, $E$, $L$), Key Step Unlearning Optimization, and Prompt Augmentation (PA) on KSCU’s performance.
[…]
KSCU early 70% steps 91.1 59.8 15.6 541
KSCU early 80% steps 92.1 58.1 14.1 624
KSCU shrink $E$ 88.5 89.3 15.3 570
KSCU last 60% steps 79.7 85.3 15.7 498
KSCU last 80% steps 96.1 48.1 14.3 628
KSCU all steps 96.0 54.5 18.8 782
KSCU $\lambda_{1}$ = 0 93.6 54.9 13.5 546
KSCU $\lambda_{2}$ = 0 92.2 57.3 14.1 546
KSCU w/o PA 92.5 51.8 14.5 549
KSCU (last 70% steps) 96.5 47.2 14.1 546
For the Key Step Unlearning Optimization, setting $\lambda_{1}=0$ disables the Unconditional Predicted Noise Optimization component. This configuration led to improved retainability but a clear decline in unlearning effectiveness. This demonstrates that unconditional noise optimization plays a central role in effective forgetting, albeit at the cost of slightly compromising generation quality. Conversely, setting $\lambda_{2}=0$ degrades the optimization objective to preserving the model’s unconditional predicted noise across training. This ablation yields noticeably weaker unlearning performance compared to full KSCU, confirming the necessity of balancing both terms. Interestingly, the generative retainability of dynamic $\lambda_{2}$ (i.e., standard KSCU) remains comparable to that of $\lambda_{2}=0$, indicating that dynamically weighting $\lambda_{1}$ effectively guides $loss_{\text{regular}}$ to preserve generation in early steps while strengthening unlearning in later ones.
Finally, for the Prompt Augmentation module, removing it resulted in a clear drop in UA and an increase in ASR, indicating reduced robustness against adversarial or paraphrased prompts. Notably, the FID remained almost unchanged, suggesting that Prompt Augmentation imposes minimal impact on generative retainability while substantially enhancing robustness. This confirms Prompt Augmentation as a lightweight yet highly effective strategy for improving resilience to real-world prompts without sacrificing image quality.
IV-F Target Key Step Concept Unlearning
In Section 3.2, we derived our loss function, which encourages the model to sample in the direction opposite to the original concept $c$. This formulation effectively prevents the generation of $c$ by pushing the model toward unrelated alternatives. However, the drawback of this approach is that the replacement concept is uncontrolled, often leading to unpredictable or semantically irrelevant outputs.
To mitigate this limitation, we introduce a slight modification to the loss function that explicitly guides the model to replace $c$ with a user-specified target concept $c^{-}$, rather than leaving the substitution unconstr ained. The modified formulation is as follows:
$\begin{split}\hat{\epsilon}^{*}_{\theta}(\mathbf{x}_{t},t,\mathbf{c})&=\epsilon_{\theta}(\mathbf{x}_{t},t)+w\cdot\underbrace{(\epsilon_{\theta}(\mathbf{x}_{t},t,\mathbf{c^{-}})-\epsilon_{\theta}(\mathbf{x}_{t},t))}_{\text{Guiding towards }c^{-}}\\ &=\epsilon_{\theta}(\mathbf{x}_{t},t)+w\cdot\underbrace{(\epsilon^{*}_{\theta}(\mathbf{x}_{t},t,\mathbf{c^{+}})-\epsilon_{\theta}(\mathbf{x}_{t},t))}_{\text{Guiding towards }c^{+}},\end{split}$ (8)
where $c^{+}$ denotes the concept to be unlearned, and $c^{-}$ is the intended replacement concept.
Building on this, we define the following modified loss function to explicitly enforce targeted substitution:
This objective provides precise control over the replacement process, ensuring that the model not only forgets $c^{+}$ but also consistently aligns its generations with the specified substitute $c^{-}$. Compared to uncontrolled forgetting, this approach transforms concept unlearning into a more structured substitution task, enabling applications such as targeted style transfer or controlled semantic replacement within diffusion models.
Figure 7 illustrates this behavior, where KSCU successfully replaces bird with balloon and Van Gogh with sketch, confirming the effectiveness of the proposed target unlearning strategy.
Figure 7: Visualization results of KSCU with target replacement. The left column shows generations from SDv1.5, while the right column illustrates the results after applying KSCU. Here, the target concepts bird and Van Gogh are replaced with balloon and sketch, respectively.
V Conclusion
This study explores the role of denoising steps in concept unlearning and proposes KSCU, which selectively targets key steps for improved effectiveness. Experiments across four tasks demonstrate that KSCU outperforms previous SOTA methods in overall performance. The key-step perspective introduced here offers valuable insights for future research.
@@ Appendix A Analysis of Frequency Component Destruction in Diffusion Models
A-A Definition of the Diffusion Process
@@
$d\mathbf{x}_{t}=g(t)d\mathbf{w}_{t},\quad\mathbf{x}_{0}\text{ as the initial condition.}$ (11)
A-B Frequency-Domain Analysis
@@
A-C Derivation of Frequency Component Degradation
A-C1 Representation of the Initial Signal and Noise
@@
A-C2 Power Spectral Density of the Signal
@@
$\mathbb{E}[|\hat{\mathbf{x}}_{t}(\omega)|^{2}]=|\hat{\mathbf{x}}_{0}(\omega)|^{2}+\int_{0}^{t}|g(s)|^{2}ds.$ (16)
A-C3 Signal-to-Noise Ratio (SNR) Analysis
@@
A-C4 Frequency Dependence of SNR Decay
@@
Since $\int_{0}^{t}|g(s)|^{2}ds$ increases with time, SNR at all frequencies declines. However, because high-frequency components have lower initial power, their SNR reaches a degradation threshold $\text{SNR}_{\text{th}}$ faster than low-frequency components. Specifically, the time at which SNR falls below $\text{SNR}_{\text{th}}$ satisfies:
$\int_{0}^{t_{\text{th}}}|g(s)|^{2}ds=\frac{|\hat{\mathbf{x}}_{0}(\omega)|^{2}}{\text{SNR}_{\text{th}}}.$ (19)
@@ Appendix B More Experimental Details
Appendix C Multi-concept Unlearning
To evaluate the effectiveness of KSCU in multi-concept unlearning, we provide visualized results 8. Our observations indicate that KSCU maintains strong performance in removing multiple concepts.